OpenStudy (anonymous):

reduce to lowest terms: (2x^3-6x^2-20x)(25-x^2)^-1

OpenStudy (lgbasallote):

change to \[\Large \frac{2x^3 - 6x^2 - 20x}{25 - x^2}\] then it gets easy :D

OpenStudy (earthcitizen):

\[(-4 x - 2 x^2)/(5 + x)\]

OpenStudy (lgbasallote):

@EarthCitizen you can still factor out 2x

OpenStudy (lgbasallote):

-2x*

OpenStudy (earthcitizen):

kk, ty

OpenStudy (anonymous):

can you show me this step by step :/

OpenStudy (lgbasallote):

factor out 2x... \[\Large \frac{2x(x^2 - 3x - 10)}{25 - x^2}\] now factor out \(25 - x^2\) \[\Large \frac{2x(x^2 - 3x - 10)}{(5 - x)(5+x)}\] now factor out \(x^2 - 3x - 10\) \[\Large \frac{2x(x-5)(x+2)}{(5-x)(5+x)}\] now we factor out -1 from 5 - x \[\Large \frac{2x\cancel{(x-5)}(x+2)}{-\cancel{(x-5)}(5+x)}\]

OpenStudy (anonymous):

thank you :)...just to clarify would simplest terms be as written above with x-5 cancelled out

OpenStudy (lgbasallote):

simply \[\large \frac{-2x(x+2)}{x+5}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you mind if i ask you another question?

OpenStudy (anonymous):

((5m-10)/(8m+4))x((3m+9)/(10m-5))x((8m^2-2)/(m^2+2m-3))